OpenStudy (anonymous):

convert into a product and simplify cos(5pi/12+theta)+cos(pi/12-theta)

OpenStudy (anonymous):

i got the answer \[-\frac{ \sqrt{2} }{ 2 }\sin (\theta)\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }\sin (-\theta)\]

OpenStudy (anonymous):

\[\cos (\frac{ 5\pi }{ 12 }+\theta)\cos(\frac{ \pi }{ 12 }-\theta)\] original question

jimthompson5910 (jim_thompson5910):

is it \[\cos (\frac{ 5\pi }{ 12 }+\theta)\cos(\frac{ \pi }{ 12 }-\theta)\] or \[\cos (\frac{ 5\pi }{ 12 }+\theta)+\cos(\frac{ \pi }{ 12 }-\theta)\]

OpenStudy (anonymous):

the second one.. sorry. mistype

jimthompson5910 (jim_thompson5910):

I'm going to use the identity \[\Large \cos(A) + \cos(B) = 2\cos\left(\frac{A+B}{2}\right)\cos\left(\frac{A-B}{2}\right)\]

OpenStudy (anonymous):

isnt it -2cos(A+B/2).... not sure if it matters.. just what our book lists.

jimthompson5910 (jim_thompson5910):

that looks like part of it

OpenStudy (anonymous):

i used the wrong identity then.. ugh.

jimthompson5910 (jim_thompson5910):

yeah the identity i wrote is correct because it is listed here http://www.sosmath.com/trig/Trig5/trig5/trig5.html

jimthompson5910 (jim_thompson5910):

in this case, A = 5pi/12 + theta B = pi/12 - theta So A+B = (5pi/12 + theta) + (pi/12 - theta) A+B = pi/2 (A+B)/2 = pi/4 ----------------------------------------- A-B = (5pi/12 + theta) - (pi/12 - theta) A-B = pi/3 + 2theta (A-B)/2 = pi/6 + theta

OpenStudy (anonymous):

so \[\frac{ \sqrt{6} }{ 2 }\cos(\theta)\]

jimthompson5910 (jim_thompson5910):

cos((A+B)/2) = cos(pi/4) = sqrt(2)/2 cos((A-B)/2) = cos(pi/6 + theta) So put this all together to get \[\Large \cos(A) + \cos(B) = 2\cos\left(\frac{A+B}{2}\right)\cos\left(\frac{A-B}{2}\right)\] \[\Large \cos\left(\frac{ 5\pi }{ 12 }+\theta\right) + \cos\left(\frac{ \pi }{ 12 }-\theta\right) = 2\left(\frac{\sqrt{2}}{2}\right)\left(\cos\left(\frac{\pi}{6} + \theta\right)\right)\] \[\Large \cos\left(\frac{ 5\pi }{ 12 }+\theta\right) + \cos\left(\frac{ \pi }{ 12 }-\theta\right) = \sqrt{2}\cos\left(\frac{\pi}{6} + \theta\right)\]

OpenStudy (anonymous):

so you dont convert the cos(pi/6) to sqrt(3)/2?

jimthompson5910 (jim_thompson5910):

no you can't do that because it's not just cos(pi/6) it's cos(pi/6 + theta)

jimthompson5910 (jim_thompson5910):

if it were just cos(pi/6), then yes, you can do that

OpenStudy (anonymous):

so just distribute the 2 to cos(pi/4) and leave the 2nd cosine term. right? just making sure.. have a quiz in an hour and a half :)

jimthompson5910 (jim_thompson5910):

yes because cos(pi/4) is really sqrt(2)/2

OpenStudy (anonymous):

perfect. thanks for your help once again. I understand now.

jimthompson5910 (jim_thompson5910):

yw